Let a function f : R -> R be defined by f(x) = {2, if x belongs to Q , 4 , if x doesn't belongs to Q} check whether f is continuous on B
1
Expert's answer
2020-03-30T06:32:29-0400

We will show that the function is not continuous at any point "x \in \mathbb{R}." Consider an arbitrary point "x \in \mathbb{R}." It is well known that the set of rational numbers "\mathbb{Q}" is everywhere dense in "\mathbb{R}." Therefore there exists a sequence "\{r_n\}_{n=1}^\infty \subset \mathbb{Q}" such that "r_n \to x" for "n \to \infty". Hence "f(r_n)=2 \to 2" for "n \to \infty". Also it is known that the set of irrational numbers "I=\mathbb{R} \setminus \mathbb{Q}" is everywhere dense in "\mathbb{R}". Therefore there exists a sequence "\{x_n\}_{n=1}^\infty \subset I" such that "x_n \to x" for "n \to \infty". Hence "f(x_n)=4 \to 4" for "n \to \infty."

So, thus we found two sequences "r_n \to x \gets x_n," but "|f(r_n)-f(x_n)| \not \to 0" for "n \to \infty". Therefore "f" is not continuous at "x \in \mathbb{R}." Now if "B \subset \mathbb{R}" is any subset, then "f" is not continuous on "B" , because "f" is discontinuous at every point.